A total of $20,000 is to be invested. Some in bonds and some in CD's. The amount in bonds is to exceed that in CD's by $3,000. How much will be invested in each?
It has been 18 years since I have done algebra problems, and I am having a little bit of difficulty. The book shows an example, but it is in a fraction form, this isn't and it is confusing me. Any help you can give me would be appreciated. Thank you

Found 2 solutions by lucky83, DWL:
Answer by lucky83(12)

the first step is to set up your equation
I am going to set x=bonds and y=cds
equation one
x+y=20000
equation 2
x=y+3000
substitute equation 2 back into equation 1
(y+3000)+y=20000
2y+3000=20000
2y=17000
y=8500(cds)
now you can take this y value and plug it back into either equation you want
x=8500+3000
x=11500(bonds)

Answer by DWL(56)

Ok We will call the bonds X and the CD's Y
We know that X + Y = 20,000
we also know that X = Y + 3,000
so lets solve the equation:
Y + 3000 + Y = 20,000
Combine terms:
2*Y+3,000=20,000
Subtract 3000 from both sides results: 2*Y=17000
Divide both sides by 2 results: Y = 8,500
so if Y = 8,500, we know X = Y + 3000, so X = 8500 + 3000 = 11,500
So bonds = 11,500
And CDs = 8,500
For a total of 20,000